### v1

@@ 3.2 Resource-efficient Finetuning
Building on the evaluation of PEFT methods, we propose resource-efficient finetuning by adopting a similar workflow to CellSeg1 [Zhou et al.(2024)Zhou, Du, and Xu], who finetune SAM for automatic segmentation on a single image using LoRA. We compare their method with our finetuning approach for automatic and interactive segmentation, using Freeze Encoder, QLoRA, LoRA, and Full Finetuning, which offer the best efficiency-quality trade-offs. For more efficiency, we lower the number of objects per image to 5, otherwise using the same settings as in Sec. 3.1. We use one image for training, one for validation, and the same test sets as before. Fig. 5 shows the results. SAM improves for all tasks, $\mu$SAM mainly for AIS. Our approach is on par with CellSeg1, for which we also use both base models, while improving interactive segmentation. Interestingly, full finetuning is on par with PEFT, despite the small training data and more trainable parameters. For large domain shifts (Platynereis) it outperforms PEFT. We also benchmark efficiency, see Tabs. 8, 9, 10. Here, we again see major gains when freezing the encoder, and small gains from other PEFT.
@@
Figure 5: Resource-efficient training for microscopy segmentation. Our methods use one image for training, one for validation, CellSeg1 uses only a single image.
@@ 4 Discussion
We conduct a thorough study of PEFT for segmentation in biomedical images with SAM. Our experiments show that PEFT achieves a similar quality as full finetuning. Contrary to observations made by others, e.g. [Dutt et al.(2024)Dutt, Ericsson, Sanchez, Tsaftaris, and Hospedales], it does not lead to better results for limited training data. The performance of PEFT methods varies across datasets, LoRA is the best overall, QLoRA is good for small domain gaps. Efficiency gains are marginal for small models (ViT-B, ViT-L) and small for larger models (ViT-H). An exception is encoder freezing, which leads to clear gains, without a strong negative impact on quality. Given these results, we recommend freezing the encoder in limited resource settings, exploring QLoRA or LoRA with medium resources, which may enable training despite only small efficiency benefits, and otherwise using full finetuning.
We propose an approach for resource-efficient finetuning, which can improve SAM with only two annotated image, similar to [Zhou et al.(2024)Zhou, Du, and Xu], but also improving interactive segmentation. We believe that this approach will speed up many practical segmentation tasks, by enabling finetuning SAM previously hindered by resource demands. To this end, we will implement our approach in the $\mu$SAM tool, which provides interactive data annotation, enabling efficient human-in-the-loop annotation and training in one tool.
Finally, we believe that our work will benefit future work on PEFT for vision foundation models, e.g. for SAM2 [Ravi et al.(2024)Ravi, Gabeur, Hu, Hu, Ryali, Ma, Khedr, Rädle, Rolland, Gustafson, et al.]. We also show a need to improve PEFT for segmentation; current methods do not yield clear advantages in the quality-efficiency trade-off. Such improvements may be achieved by simple strategies, such as using adapters in a subset of layers, or through more fundamental advances.
@@
The work of Anwai Archit was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - PA 4341/2-1. This work was also supported by the Google Research Scholarship “Vision Foundation Models for Bioimage Segmentation”. We gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy at NHR@Göttingen as part of the NHR infrastructure, under the project nim00007. We acknowledge the use of icons from Bioicons [Jug and Contributors(2023)] in Fig. LABEL:fig:peft_sam. In particular, the microscope icon by DBCLS [for Life Science (DBCLS)(2023)] is licensed under CC-BY 4.0.
Appendix A Evaluation Metric
@@
Appendix B PEFT Evaluation
B.1 Segmentation Quality
@@
Fig. 3 and Fig. 4 present a comparison of various PEFT methods across microscopy and medical datasets respectively. The PEFT methods are arranged from left to right in the order of the number of trainable parameters, with full finetuning having the most parameters. Table 5 lists the number of parameters for each method. We evaluate the methods using training from the default SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] and from the generalist $\mu$-SAM models [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape], where we use either the electron microscopy or light microscopy model, and from MedicoSAM [Archit et al.(2025)Archit, Freckmann, and Pape] for medical imaging. The models are evaluated on automatic instance segmentation (AIS) for microscopy experiments, and initial prompts (point and box prompts), and iterative prompting, beginning either from a box prompt ($I_{B}$) or a point prompt ($I_{P}$) for both microscopy and medical imaging experiments.
@@
As an overall trend, we observe that as the number of trainable parameters increases, the segmentation quality also tends to increase across the different datasets, demonstrating a tradeoff between model complexity, capacity, and performance. The relative performance of the PEFT methods varies between datasets, indicating that the suitability of different PEFT methods may depend on the dataset’s characteristics, such as complexity, size, and the model’s initial performance. Despite this variation, full finetuning, Attention Tuning, and LoRA consistently show good performance; QLoRA shows good performance for the domain specific foundation models ($\mu$SAM, MedicoSAM), but performs bad for default SAM, which has a larger domain gap. Among the listed PEFT methods, Attention Tuning has the highest number of trainable parameters, requiring 32.5M parameters. In contrast, LoRA significantly reduces the parameter space to just 5.4M parameters, amounting to only 5.6% of the parameters used in Full Finetuning (FFT). Despite this substantial reduction, LoRA achieves performance that is on par with Attention Tuning and Full Finetuning in most cases. To simplify these observations, we averaged the performance of the different PEFT methods over all datasets in Tabs. 1 (microscopy, SAM), 2 (microscopy, $\mu$SAM), 3 (medical imaging, SAM), and 4 (medical imaging, MedicoSAM). LoRA consistently ranks among the top three methods, frequently alongside full fine-tuning and attention tuning, despite these methods having significantly more parameters. It is notable that for the iterative prompting task, FacT performs very well alongside LoRA. This is especially interesting because FacT is essentially an extension of LoRA, applying the low-rank decomposition to a tensorized vision transformer. Moreover, while some PEFT methods excel for specific tasks - such as Bias Tuning for Mitolab - LoRA consistently ranks among the top methods across diverse scenarios, including AIS and single point and box prompt. This makes LoRA the best PEFT choice from a segmentation quality perspective.
The strong performance of both LoRA and Attention Tuning suggests that tuning the attention weight matrices is particularly important for domain adaptation. This highlights the potential benefit of focusing on attention-based techniques when fine-tuning models for specialized tasks.
@@
Full FT **0.406 *0.351 **0.696 0.793 0.849
Attn Tune *0.414 **0.337 ***0.695 **0.803 **0.852
AdaptFormer 0.334 0.287 0.672 0.788 0.837
LoRA **0.390 ***0.336 *0.702 *0.804 ***0.852
@@
Table 3: The dice coefficient of different PEFT methods, trained from default SAM, for medical segmentation and averaged over all 6 datasets. (*): 1st Place; (**): 2nd Place; (***): 3rd Place.
@@
Full FT *0.723 **0.830 0.811 0.850
Attn Tune **0.721 *0.810 **0.825 0.855
AdaptFormer 0.612 0.814 0.789 ***0.858
LoRA ***0.673 ***0.826 *0.838 *0.862
QLoRA 0.386 0.757 0.394 0.508
FacT 0.633 0.812 ***0.815 **0.862
SSF 0.592 0.802 0.803 0.850
Bias Tune 0.615 0.810 0.802 0.847
LN Tune 0.596 0.818 0.805 0.843
Freeze Encoder 0.521 0.791 0.784 0.833
Table 4: The dice coefficient of different PEFT methods, trained from MedicoSAM, for medical segmentation and averaged over all 6 datasets. (*): 1st Place; (**): 2nd Place; (***): 3rd Place.
@@
Full FT *0.709 0.813 0.826 0.868
Attn Tune ***0.687 0.824 **0.862 0.879
AdaptFormer 0.670 0.842 0.856 0.880
LoRA **0.694 0.836 *0.883 *0.901
QLoRA 0.639 ***0.846 ***0.856 **0.890
FacT 0.645 **0.849 0.852 0.885
SSF 0.658 0.845 0.853 0.883
Bias Tune 0.643 0.832 0.845 0.881
LN Tune 0.645 *0.850 0.853 0.879
Freeze Encoder 0.583 0.812 0.830 0.879
@@
fig:qualitative_iterative_prompting Refer to caption
@@
fig:qualitative_iterative_prompting Refer to caption
@@
B.2 Computational Efficiency
We measure the efficiency of PEFT methods, in terms of number of trainable parameters, training time per iteration and in total, and VRAM memory usage during training. Tab. 5 reports these measures for ViT-B and all PEFT methods, Tab. 6 and Tab. 7 report them for ViT-L and ViT-H respectively, for only a subset of PEFT methods. These results were obtained from finetuning default SAM on the LIVECell [Edlund et al.(2021)Edlund, Jackson, Khalid, Bevan, Dale, Dengel, Ahmed, Trygg, and Sjögren] dataset, a large dataset with annotations for cell segmentation in phase-contrast microscopy. Note that we perform all experiments for segmentation quality with the smallest model, ViT-B, because it was shown e.g. in [Gu et al.(2024)Gu, Dong, Yang, and Mazurowski, Archit et al.(2025)Archit, Freckmann, and Pape] that using it does not have a noticeable impact compared to the larger models ViT-L and ViT-H for segmentation quality in biomedical imaging. However, we evaluate the computational efficiency for all model sizes, as this observation may not hold true for other domains, for which finetuning the larger models may be beneficial. We report efficiencies only for a single dataset, as these measurements are largely independent of the data specifics. The exception is the overall training time because we use early stopping. Specifically, we stop the training after 10 epochs without an improvement in the validation score. Hence, the overall training time depends on dataset characteristics. We report the training time across datasets in Fig. 8.
We see that all PEFT methods lead to a clear reduction in the number of trainable parameters, from 30% (Attn Tune) to less than ca. 5% (all others). However, this reduction in trainable parameters does not result in a clear reduction of memory requirements or training time per iteration. This fact is especially noticeable for ViT-B, for the larger models, especially ViT-H, the reduction in memory becomes a bit more pronounced. The exception is freezing the encoder, which clearly reduces the memory demand and also reduces training time per iteration, for all model sizes. It is unclear to us why the reduction in trainable parameters is not accompanied by a similar reduction in memory consumption for other methods.

### v2

@@ 3.2 Resource-efficient Finetuning
Building on the evaluation of PEFT methods, we propose resource-efficient finetuning by adopting a similar workflow to CellSeg1 [Zhou et al.(2024)Zhou, Du, and Xu], who finetune SAM for automatic segmentation on a single image using LoRA. We compare their method with our finetuning approach, explained in detail in App. C for automatic and interactive segmentation, using Freeze Encoder, QLoRA, LoRA (including late PEFT settings for all 3), and Full Finetuning, which offer the best efficiency-quality trade-offs. To reduce hardware demand, we lower the number of objects per image to 5, otherwise using the same settings as in Sec. 3.1. We use one image for training, one for validation, and the same test sets as before. Fig. 4 shows the results for microscopy and medical data (only interactive segmentation for the latter). SAM improves for all tasks, $\mu$SAM mainly for AIS. Our approach is on par with CellSeg1, for which we also use both base models, while also improving interactive segmentation. Interestingly, full finetuning is on par with PEFT, despite the small training data and having more trainable parameters. For large domain shifts (Platynereis) it outperforms PEFT. We also benchmark efficiency, see Tabs. 8, 9, 10. Here, we again see major gains when freezing the encoder, small gains from other PEFT, and a good trade-off using “late PEFT”. Further, we investigate the effect of adding more training images in Fig. 12.
@@
Figure 4: Resource-efficient training for microscopy and medical segmentation. We compare several PEFT methods in data limited settings. Our methods use one image for training, one for validation, CellSeg1 uses only a single image.
@@ 4 Discussion
We conducted a thorough study of PEFT for segmentation in biomedical images with SAM. Our experiments show that PEFT achieves a similar quality as full finetuning. Contrary to observations made by others, e.g. [Dutt et al.(2024)Dutt, Ericsson, Sanchez, Tsaftaris, and Hospedales], it does not lead to better results for limited training data compared to full FT. The performance of PEFT methods varies across datasets, LoRA is the best overall, QLoRA is good for small domain gaps. Efficiency gains are marginal for standard PEFT, with the exception of encoder freezing, due to the fact that our tasks are activation rather than parameter bound. Based on these findings we introduce late PEFT, which improves efficiency with only small loss of quality. Given these results, we recommend freezing the encoder in very limited resource settings (e.g. on a CPU), using late PEFT for limited resources (e.g. consumer GPU), and otherwise using full finetuning (high-end GPU). See App.  E.3 for more detailed recommendations.
We proposed an approach for resource-efficient finetuning, which can improve SAM with only two annotated image, similar to [Zhou et al.(2024)Zhou, Du, and Xu], but also improves interactive segmentation. We believe that this approach will speed up many practical segmentation tasks, by enabling finetuning SAM previously hindered by resource demands. We have integrated this approach in the $\mu$SAM tool, which provides interactive data annotation, enabling efficient human-in-the-loop annotation and training in one tool.
We believe that our work will benefit future work on PEFT for (vision) foundation models, e.g. for SAM2 [Ravi et al.(2025)Ravi, Gabeur, Hu, Hu, Ryali, Ma, Khedr, Rädle, Rolland, Gustafson, et al.]. Our results on late PEFT show how such strategies need to be adapted for vision transformers, which are activation rather than parameter bound, offering a clear direction for future research.
@@
fig:late_lora Refer to caption
Figure 5: a) Parameter count vs. memory requirements during training for Late LoRA, Late QLoRA and late finetuning (Late FT) for microscopy finetuning on HPA. We report two different LoRA approaches, where adapters are applied only to the attention layer (LoRA-C) or also to the MLP (LoRA-A). b) Same as a., but for finetuning on medical data (PSFHS). Note that the overall memory requirement is significantly lower because we do not train the encoder for automatic segmentation, which requires an additional forward and backward pass during training. c) Segmentation quality as a function of late finetuning on HPA and PSFHS.
[…]
@@
The work of Anwai Archit was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - PA 4341/2-1. Constantin Pape is supported by the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) under under Germany’s Excellence Strategy - EXC 2067/1-390729940. This work is supported by the Ministry of Science and Culture of Lower Saxony through funds from the program zukunft.niedersachsen of the Volkswagen Foundation for the ’CAIMed – Lower Saxony Center for Artificial Intelligence and Causal Methods in Medicine’ project (grant no. ZN4257). It was also supported by the Google Research Scholarship “Vision Foundation Models for Bioimage Segmentation”. We gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy at NHR@Göttingen as part of the NHR infrastructure, under the project nim00007. We acknowledge the use of icons from Bioicons [Jug and Contributors(2023)] in Fig. 1. In particular, the microscope icon by DBCLS [for Life Science (DBCLS)(2023)] is licensed under CC-BY 4.0.
Appendix A Efficiency of PEFT
For the optimal use of PEFT methods in vision transformers, it is important to understand the main contributors to the memory footprint during training. Unlike typical LLM training, where trainable parameters dominate memory usage, training SAM (and likely other ViT applications) is primarily constrained by activation gradients. The distinction becomes evident if we consider the ratio of parameters and sequence length, which controls the size of activations. To illustrate this, consider the following: GPT-3, with a sequence length of 2048 tokens and 175B parameters [Brown et al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan, Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess, Clark, Berner, McCandlish, Radford, Sutskever, and Amodei], yields a ratio of approximately 85.5M. LLaMa-2 [Touvron et al.(2023)Touvron, Martin, Stone, Albert, Almahairi, Babaei, Bashlykov, Batra, Bhargava, Bhosale, Bikel, Blecher, Ferrer, Chen, Cucurull, Esiobu, Fernandes, Fu, Fu, Fuller, Gao, Goswami, Goyal, Hartshorn, Hosseini, Hou, Inan, Kardas, Kerkez, Khabsa, Kloumann, Korenev, Koura, Lachaux, Lavril, Lee, Liskovich, Lu, Mao, Martinet, Mihaylov, Mishra, Molybog, Nie, Poulton, Reizenstein, Rungta, Saladi, Schelten, Silva, Smith, Subramanian, Tan, Tang, Taylor, Williams, Kuan, Xu, Yan, Zarov, Zhang, Fan, Kambadur, Narang, Rodriguez, Stojnic, Edunov, and Scialom], with a sequence length of 4096 and 70B parameters has a ratio of 17M. BERT-base, used with a sequence length of 512 and 110M parameters [Devlin et al.(2019)Devlin, Chang, Lee, and Toutanova], results in a ratio of 215,000. In contrast, for the Segment Anything Model (SAM), an input image size of 1024x1024 [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick], divided into 16x16 patches, results in a sequence length of 64x64 (4,096). For ViT-B, with 86M parameters [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby], this corresponds to a parameter-to-sequence length ratio of approximately 20,000.
While methods like LoRA effectively reduce the number of trainable parameters, they do not address activation gradients directly. Meaningful memory savings only occur when the initial layers of the model are frozen, allowing the optimizer to discard the corresponding activation gradients.
To test this hypothesis, we conducted an experiment using SAM trained on LIVECell [Edlund et al.(2021)Edlund, Jackson, Khalid, Bevan, Dale, Dengel, Ahmed, Trygg, and Sjögren] data with a batch size of 1. We froze the entire image encoder except for a single block and observed the following results: unfreezing the first block of the image encoder reduced memory usage by approximately 500MB, while unfreezing the last block saved around 5GB compared to full fine-tuning.
These findings motivated the concept of Late PEFT. For the example in Late LoRA, LoRA adapters are applied only from a certain point onward in the image encoder, reducing memory usage by limiting activation gradient calculations, while still providing updates to the image encoder.
[Dettmers et al.(2023)Dettmers, Pagnoni, Holtzman, and Zettlemoyer] suggest that once memory constraints are taken care of, performance can be further optimized by introducing additional trainable parameters through more low-rank adapters. Rather than limiting LoRA to the query and value matrices of the attention blocks, which is a common practice recommended by the original LoRA authors [Hu et al.(2022)Hu, Shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen], we explore the application of LoRA to all linear layers within the attention blocks. We distinguish these approaches as “LoRA Classic” (LoRA-C) and “LoRA All” (LoRA-A) in Fig. 5.
We explore late PEFT methods by applying PEFT only from a specific attention block onward in the encoder, avoiding backpropagating gradients through the entire encoder, thereby reducing memory usage. We evaluate this approach using LoRA, QLoRA and full fine-tuning, applying each method to 50%, 25%, and 8% of the attention blocks. Our results in Fig. LABEL:fig:lateft_microscopy and LABEL:fig:lateft_medical show that the latter approach, applied to 50% of the blocks achieves the best trade-off between performance and memory efficiency.
Appendix B Implementation of QLoRA
For the implementation of QLoRA, we make use of bitsandbytes (https://huggingface.co/docs/bitsandbytes/main) for quantizing the linear layers. The implementation expects users to indicate linear layers to map to a lower precision (we reduce the precision down to 4 bit and freeze them, as training for this precision is a technical limitation in PyTorch). This necessitates converting the pretrained weights to the target precision (4 bit), initializing the quantized layers accordingly, and finally injecting LoRA to update the attention matrices. During training, the quantized layers are frozen, i.e. not updated through gradients. Their primary task is to improve efficiency; the LoRA matrices are learned at full precision. Another important finding is to avoid compound operations with tensors which require gradients, otherwise the PyTorch optimizer cannot trace gradients due to quantization in the computation graph, a known limitation of in-place operations in PyTorch. This contributed to the challenge of setting up a proper training implementation with quantization and mixed precision, combined with PyTorch optimizer specifics, in such a sensitive setup.
In inference, the approach to use the finetunined QLoRA model with quantized 4-bit layers, except for LoRA parameters of the attention matrices, did not work. After some empirical investigation, our analysis led us to introduce a new setup, where inferring with QLoRA model expects all layers to be initialized with the pretrained weights in the original precision, and replace the attention matrices with the learned LoRA weights. This resulted in meaningful improvements and comparable results to other PEFT methods in settings with small domain shifts.
Appendix C Implementation of Resource-efficient Finetuning
For microscopy, the choice of the 2 images for training was done manually by selecting images with a good representation of the overall task. CellSeg1 [Zhou et al.(2024)Zhou, Du, and Xu] has shown that the choice of image is crucial in this setting. Once these two images are selected, we use 1 image for training and the other exclusively for validation, and test our trained model on the entire test set, for consistency with other experiments. The same goes for medical imaging, however for 3D images, e.g. MOTUM and Mice TumSeg, we choose one plane with the foreground object annotations for training, and another from a different volume for validation. This step in particular is critical to avoid any negative bias, e.g. selecting an image with underrepresented / sparsely distributed cells or a 3D plane with an odd morphology of an organ.
We use CellSeg1 for benchmarking purpose with our data efficient experiments. Unlike CellSeg1, we improve SAM both for interactive and automatic segmentation. CellSeg1 only improves automatic segmentation, and is thus also not applicable for our medical experiments. Furthermore, we offer easy-to-use example scripts and a command line interface for finetuning. Our models are compatible with $\mu$SAM and the finetuning strategies without PEFT are directly compatible in several other annotation softwares, which is a clear limitation of CellSeg1 that can only be used within their tool.
Appendix D Evaluation Metric
@@
Appendix E PEFT Evaluation
E.1 Segmentation Quality
@@
Fig. 3 presents a comparison of various PEFT methods across microscopy and medical datasets respectively. The PEFT methods are arranged from left to right in the order of the number of trainable parameters, with full finetuning having the most parameters. Table 5 lists the number of parameters for each method. We evaluate the methods using training from the default SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick] and from the generalist $\mu$SAM models [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape], where we use either the electron microscopy or light microscopy model, and from MedicoSAM [Archit et al.(2025b)Archit, Freckmann, and Pape] for medical imaging. The models are evaluated on automatic instance segmentation (AIS) for microscopy experiments, and point and box prompts for both microscopy and medical imaging experiments. We also report a simulated interactive segmentation setup, which we call the iterative prompting task, i.e. starting with either a single point ($I_{P}$) or a box prompt ($I_{B}$) for each object, and iterative rectification with additional points to improve the segmentation quality
@@
As an overall trend, we observe that as the number of trainable parameters increases, the segmentation quality also tends to increase across the different datasets, demonstrating a trade-off between model complexity, capacity, and performance. The relative performance of the PEFT methods varies between datasets, indicating that the suitability of different PEFT methods may depend on the dataset’s characteristics, such as complexity, size, and the model’s initial performance. Despite this variation, Full Finetuning, Attention Tuning, and LoRA consistently show good performance. Two of the late PEFT methods, Late LoRA and Late FT, show some reduction in performance, but with relative minor differences to their “non-late” counterparts. (Late) QLoRA shows good performance for the domain specific foundation models ($\mu$SAM, MedicoSAM), but performs bad for default SAM, which has a larger domain gap.
To simplify these observations, we averaged the performance of the different PEFT methods over all datasets in Tabs. 1 (microscopy, SAM), 2 (microscopy, $\mu$SAM), 3 (medical imaging, SAM), and 4 (medical imaging, MedicoSAM). These averaged results validate our observations from above, highlighting that Attention Tuning, (Late) LoRA and Late FT show consistently good improvements. While some PEFT methods excel for specific tasks - such as Bias Tuning for MitoLab - they are not competitive with these other methods.
@@
Late FT 0.336 ***0.300 0.615 0.675 0.753
Late LoRA 0.294 0.275 0.632 0.693 0.764
Late QLoRA 0.297 0.203 0.562 0.197 0.299
AdaptFormer 0.334 0.287 0.672 0.788 0.837
@@
Full FT **0.406 *0.351 ***0.696 0.793 0.849
Attn Tune *0.414 **0.337 **0.695 **0.803 **0.852
Late FT 0.371 0.315 0.639 0.752 0.829
Late LoRA ***0.393 0.333 0.675 0.790 0.844
Late QLoRA 0.334 0.295 0.621 0.746 0.807
LoRA 0.390 ***0.336 *0.702 *0.804 ***0.852
@@
Table 3: The dice coefficient of different PEFT methods, trained from default SAM, for medical segmentation and averaged over all 9 datasets. (*): 1st Place; (**): 2nd Place; (***): 3rd Place.
@@
Full FT **0.511 ***0.685 0.709 0.779
Attn Tune *0.522 *0.647 *0.709 **0.779
Late FT 0.405 0.687 0.685 0.756
Late LoRA 0.391 **0.709 ***0.717 *0.797
Late QLoRA 0.288 0.560 0.169 0.272
AdaptFormer 0.335 0.648 0.678 0.753
LoRA ***0.424 0.694 **0.726 ***0.788
QLoRA 0.296 0.518 0.218 0.286
FacT 0.373 0.667 0.703 0.765
SSF 0.302 0.636 0.669 0.754
Bias Tune 0.367 0.647 0.710 0.772
LN Tune 0.335 0.653 0.684 0.749
Freeze Encoder 0.255 0.581 0.607 0.705
Table 4: The dice coefficient of different PEFT methods, trained from MedicoSAM, for medical segmentation and averaged over all 9 datasets. (*): 1st Place; (**): 2nd Place; (***): 3rd Place.
@@
Full FT **0.513 0.717 0.772 0.835
Attn Tune *0.516 *0.733 *0.815 ***0.858
Late FT 0.406 **0.731 0.803 **0.859
Late LoRA 0.365 0.723 0.763 0.847
Late QLoRA 0.389 0.707 0.798 0.851
AdaptFormer 0.443 0.707 ***0.805 0.842
LoRA ***0.485 ***0.730 **0.812 *0.863
QLoRA 0.422 0.708 0.797 0.854
FacT 0.445 0.709 0.802 0.855
SSF 0.408 0.699 0.790 0.843
Bias Tune 0.414 0.692 0.785 0.841
LN Tune 0.409 0.697 0.781 0.835
Freeze Encoder 0.345 0.653 0.740 0.822
@@
fig:quanti_iterative_prompting_microscopy Refer to caption
Figure 6: Results for iterative prompting on 6 different microscopy datasets, starting with point or box prompts. The accuracy is evaluated after 8 iterations of predicting and correcting prompts.
[…]
fig:quanti_iterative_prompting_medical Refer to caption
Figure 7: Results for iterative prompting on 9 different medical imaging datasets, starting with point or box prompts. The accuracy is evaluated after 8 iterations of predicting and correcting prompts. The late finetuning experiments are only run on the core 6 datasets.
[…]
fig:qualitative_iterative_prompting_microscopy Refer to caption
@@
fig:qualitative_iterative_prompting_medical Refer to caption
@@
E.2 Computational Efficiency
We measure the efficiency of PEFT methods, in terms of number of trainable parameters, training time per iteration and in total, and VRAM memory usage during training. Tab. 5 reports these measures for ViT-B and all PEFT methods, Tab. 6 and Tab. 7 report them for ViT-L and ViT-H respectively, for only a subset of PEFT methods. These results were obtained from finetuning default SAM on the LIVECell [Edlund et al.(2021)Edlund, Jackson, Khalid, Bevan, Dale, Dengel, Ahmed, Trygg, and Sjögren] dataset, a large dataset with annotations for cell segmentation in phase-contrast microscopy. Note that we perform all experiments for segmentation quality with the smallest model, ViT-B, because it was shown e.g. in [Gu et al.(2025)Gu, Dong, Yang, and Mazurowski, Archit et al.(2025b)Archit, Freckmann, and Pape] that using it does not have a noticeable impact compared to the larger models ViT-L and ViT-H for segmentation quality in biomedical imaging. However, we evaluate the computational efficiency for all model sizes, as this observation may not hold true for other domains, for which finetuning the larger models may be beneficial. We report efficiencies only for a single dataset, as these measurements are largely independent of the data specifics. The exception is the overall training time because we use early stopping. Specifically, we stop the training after 10 epochs without an improvement in the validation score. Hence, the overall training time depends on dataset characteristics. We report the training time across datasets in Fig. 10 and report the average over all datasets for each PEFT method in the same figure.
We see that all PEFT methods lead to a clear reduction in the number of trainable parameters, from 30% (Attn Tune) to less than ca. 5% (all others). However, this reduction in trainable parameters does not result in a clear reduction of memory requirements or training time per iteration. This fact is especially noticeable for ViT-B, for the larger models, especially ViT-H, the reduction in memory becomes a bit more pronounced. The exception is freezing the encoder, which clearly reduces the memory demand and also reduces training time per iteration, for all model sizes. With our new findings reported in Appendix A for Late PEFT, a future direction is quite clear for increasing efficiency of parameter efficient finetuning methods.